OpenStudy (anonymous):

how do i find the slope in 2x-3y=9

OpenStudy (anonymous):

convert it to y = mx + b form, then m will be slope

OpenStudy (anonymous):

You move 2x over to the other side of the equation so it's -3y=-2x+9 and then divide each side by -3 so it's y= 2/3x-3

OpenStudy (anonymous):

2x-3y=9 2x - 9 = 3y 2 9 --x - -- = y 3 3 2 --x - 3 = y 3 so slope = 2/3

OpenStudy (amistre64):

define to points and work the slope between them

OpenStudy (amistre64):

when x=0, y = -3 when x=1; y= -7/3 the change in y is therefore -7/3 + 3 = 2/3

OpenStudy (anonymous):

There's an easier way than to convert it into y= mx + b. It's currently in A + B = C form. A = 2 B = -3 C = 9 To find the slope, all you do is negative A over B So -2/-3 = 2/3 Slope is 2/3